40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 8 18:35:09 2019, gautam, Update, LSC, Locking prep 
    Reply  Wed Oct 9 18:15:26 2019, gautam, Update, LSC, PRMI 3f locking PRMI_1f.png
       Reply  Wed Oct 9 22:02:58 2019, gautam, Update, LSC, REFL55 whitening issue 
          Reply  Thu Oct 10 01:12:56 2019, gautam, Update, LSC, Locking studies PRFPMIstudies.png
             Reply  Wed Oct 16 11:42:17 2019, gautam, Update, LSC, Poor separation of PRCL/MICH in 3f signals sensMat.pdf
                Reply  Thu Oct 17 11:19:28 2019, gautam, Update, LSC, Locking activity last night MICH2PRCLnulling.pdfconsolidatedSensingMatrices.pdf.zip
                   Reply  Tue Oct 22 00:52:27 2019, gautam, Update, LSC, Locking updates PRMI3f_ALS_Oct21sensMat.pdf
                      Reply  Tue Oct 29 01:24:45 2019, gautam, Update, LSC, More locking updates POP22_feature.pngPRMI3f_ALS_Oct21sensMat.pdf
                         Reply  Tue Oct 29 15:13:19 2019, gautam, Update, LSC, More locking updates PRMI_lockloss.png
                            Reply  Wed Oct 30 01:27:00 2019, gautam, Update, LSC, More locking updates POP_ASS.png
                               Reply  Wed Oct 30 11:53:41 2019, gautam, Update, LSC, MICH loop shape tuning MICH_ELP.pdfMICH_ELP_TFs.pdf
                               Reply  Wed Oct 30 17:08:40 2019, gautam, Update, LSC, POP22 investigation POP22anomaly.pdf
                                  Reply  Mon Nov 4 15:29:47 2019, gautam, Update, LSC, POP signal path POPchain.pdf
                                     Reply  Mon Nov 4 16:06:58 2019, gautam, Update, LSC, POP optical path IMG_8027.JPGlensRealignment.jpgPOPrealigned.pngPOPdither.pngPRMfixed.png
                                        Reply  Tue Nov 5 11:52:27 2019, gautam, Update, LSC, Locking notes PRMI3f_ALS_Nov4sensMat.pdf
                                           Reply  Wed Nov 6 02:08:48 2019, gautam, Update, LSC, Locking updates 
                                              Reply  Wed Nov 6 17:05:45 2019, gautam, Update, LSC, CARM calibration CARMscan.pdfCARMcalib.pdfscan.hdf5.zip
                                              Reply  Wed Nov 6 17:45:34 2019, gautam, Update, LSC, ~ 
                                                 Reply  Mon Nov 18 21:04:38 2019, gautam, Update, LSC, Locking - some ideas 
                                                    Reply  Wed Nov 20 01:07:18 2019, gautam, Update, LSC, Locking - progress PRFPMI.png
                                                       Reply  Wed Nov 20 12:14:17 2019, gautam, Update, LSC, Locking - progress TRX_TRY_comparison.pdf
                                                          Reply  Wed Nov 20 17:20:24 2019, Yehonathan, Update, LSC, QPD Investigation 
                                                             Reply  Wed Nov 20 17:52:00 2019, gautam, Update, LSC, QPD MEDM screen update newLookQPD.pngTRX_TRY_comparison.pdfIMG_8186.JPG
                                                                Reply  Wed Nov 20 21:29:28 2019, gautam, Update, LSC, PRG ~13 PRG13.pdf
                                                                   Reply  Fri Nov 22 00:54:14 2019, gautam, Update, LSC, locking notes ALSnoiseIncrease.pdf
                                                                      Reply  Mon Nov 25 19:11:22 2019, gautam, Update, LSC, ALS noise re-look ALSnoise.pdf
                                                                         Reply  Wed Nov 27 23:24:01 2019, gautam, Update, LSC, No shaking but no inspiration either 
                                                                            Reply  Sun Dec 1 13:38:48 2019, rana, Update, LSC, No shaking but no inspiration either 
                                                                               Reply  Tue Dec 3 00:03:57 2019, gautam, Update, LSC, Green ALS also shows elevated noise with high arm buildup ALSnoiseIncrease_greenBeat.pdf
                                                                                  Reply  Wed Dec 4 19:54:27 2019, gautam, Update, LSC, A look to the past DRFPMI_2016March.pdf
                                                                                     Reply  Thu Dec 5 01:54:39 2019, gautam, Update, LSC, Partial CM board path engaged AOengaged.pdf
                                                                                        Reply  Fri Dec 6 00:02:48 2019, gautam, Update, LSC, What is the correct way to set the 3f offsets? 
                                     Reply  Fri Nov 15 12:08:04 2019, gautam, Update, LSC, POPDC whitening board 
                         Reply  Tue Oct 29 17:40:48 2019, gautam, Update, LSC, More locking updates broadSweep.pdfzoomSweep.pdf
Message ID: 15062     Entry time: Tue Dec 3 00:03:57 2019     In reply to: 15057     Reply to this: 15073
Author: gautam 
Type: Update 
Category: LSC 
Subject: Green ALS also shows elevated noise with high arm buildup 

Summary:

  1. While noisier, I was able to control the arm lengths to ~30pm RMS(!) using the green ALS beats as error signals (cf. ~10 pm RMS with the IR ALS system).
  2. The PRMI could be locked with a CARM offset applied.
  3. When lowering the CARM offset, I saw an increase in the in-loop ALS error signal, just as I had with the IR beat.
  4. IR TRX / TRY unsurprisingly did not stabilize in any meaningful way.cool
  5. The noise increase seems to have some periodicity along the frequency axis - need to think about what this means.
  6. Since there is no apparent benefit to using the green ALS beats, I restored the IR system. The green PDs should still retain somewhat good alignment if one wishes to do a comparison measurement.
  7. While the shadow sensors of the ITMs report elevated noise, it is unlikely to be responsible for the cavity moving by the amount suggested by the elevated ALS error signals because of the digital low-pass filtering and 1/f^2 of the pendulum.
  8. I confirmed that the ITM shadow sensors do not report elevated noise when the PRMI is locked such that the carrier is resonant. In this config, there is comparable circulating power in the PRC as to when the CARM offset is reduced to ~0.
  9. The fact that the IR and green beats both show similar increase in noise suggestes that the cavity length / laser frequency is in fact being modulated, but I still don't know what the exact mechanism is.

was worth a shot i guess.

Trawling through some elogs, I see that this kind of feature showing up in the ALS CARM is not a new problem, see for example here. But I can't find out what the resolution was.

Attachment 1: ALSnoiseIncrease_greenBeat.pdf  24 kB  | Show | Show all
ELOG V3.1.3-